You deposit $2000 in an account that pays 2% annual interest compounded quarterly. How long will it take for the balance to reach $2400?
How do you set up the equation and rules for solving please

Answer by stanbon(75887)

You deposit $2000 in an account that pays 2% annual interest compounded quarterly. How long will it take for the balance to reach $2400?
-----------
A(t) = P(1+(r/n))^(nt)
----
2400 = 2000(1+(0.02/2))^(2t)
----
1.2 = (1+0.01)^(2t)
Take the log:
(2t)log(1.01) = log(1.2)
2t = 18.32
t = 9.165 years
-----
Cheers,
Stan H.